# Thinking in Space: How Multimodal Large Language Models See, Remember, and Recall Spaces

### v1

@@ 1 Introduction
Visual-spatial intelligence entails perceiving and mentally manipulating spatial relationships [26]; it requires myriad capabilities, including relational reasoning and the ability to transform between egocentric and allocentric perspectives (Sec.˜2). While Large Language Models (LLMs) [59, 85, 6, 100, 35, 65, 66, 9, 79, 80, 3, 75] have advanced linguistic intelligence, visual-spatial intelligence remains under-explored, despite its relevance to robotics [21, 8, 7, 62], autonomous driving [77], and AR/VR [12, 27, 53].
Multimodal Large Language Models (MLLMs) [33, 76, 47, 1, 41, 47, 4, 15], which integrate language and vision, exhibit strong capacities to think and reason in open-ended dialog and practical tasks like web agents [34, 28, 32, 21]. To advance this intelligence in the visual-spatial realm, we introduce VSI-Bench, a video-based benchmark featuring over 5,000 question-answer pairs across nearly 290 real indoor-scene videos (Sec.˜3). Video data, by capturing continuous, temporal input, both parallels how we observe the world and enables richer spatial understanding and reasoning than static images. Evaluating open- and closed-source models on VSI-Bench reveals that even though a large performance gap exists between models and humans, MLLMs exhibit emerging visual-spatial intelligence despite the challenges of video understanding, textual understanding, and spatial reasoning (Sec.˜4).
To analyze model behavior and inspired by dual-coding theory [18], which posits that linguistic and visual processing are distinct yet complementary, we prompt selected models for self-explanations (linguistic) and cognitive maps (visual). Analyzing the self-explanations reveals that spatial reasoning, as compared to visual perception, linguistic intelligence, or temporal processing, is the main factor behind weak performance on VSI-Bench (Sec.˜5). “Cognitive maps”, which represent internal layouts of environments [78, 60], allow us to evaluate MLLMs’ implicit spatial world models and find that MLLMs build strong local models but weak global ones (Sec.˜6). Furthermore, standard linguistic reasoning techniques fail to enhance performance on our benchmark. However, explicitly generating and using cognitive maps improves spatial distance question-answering.
@@ 2 Visual-Spatial Intelligence
Taxonomy. We provide a taxonomy of capabilities potentially required for visual-spatial intelligence (Fig.˜2), based on cognitive psychology [26, 55, 60, 11] and human experience with our benchmark tasks in Sec.˜3. Visual perception, linguistic intelligence, temporal processing, and spatial reasoning are the four areas needed in VSI-Bench. For example, [11] shows that visual object and spatial processing are neurally distinct, which motivates “visual perception” and “spatial reasoning” as separate areas. We break spatial reasoning into two broad capabilities: relational reasoning and egocentric-allocentric transformation.
@@ 3.1 Overview
VSI-Bench includes eight tasks of three types: configurational, measurement estimation, and spatiotemporal. The configurational tasks (object count, relative distance, relative direction, route plan) test a model’s understanding of the configuration of a space and are more intuitive for humans (see Sec.˜4 for comparison between MLLM and human performance). Measurement estimation (of object size, room size, and absolute distance) is of value to any embodied agent. While predicting a measurement exactly is very difficult, for both humans and models, a better sense of distance and other measurements is intuitively correlated with better visual-spatial intelligence and underpins a wide range of tasks that require spatial awareness, like interaction with objects and navigation. Spatiotemporal tasks like appearance order test a model’s memory of a space as seen in video. See Fig.˜3 for an overview of VSI-Bench tasks and Fig.˜5 for dataset statistics.
@@ 3.2 Benchmark Construction
We develop a sophisticated benchmark construction pipeline to effectively generate high-quality question-answer (QA) pairs at scale, as shown in Fig.˜4.
@@
Question-Answer Generation. QA pairs are primarily auto-annotated using the meta-information and question templates; the route plan task was human-annotated. We sophisticatedly design and refine the question template for each task and provide guidelines for human annotators. For more detailed design, see Sec.˜B.1.
@@
InternVL2-2B 11 27.4 21.8 24.9 22.0 35.0 33.8 44.2 30.5 7.1
InternVL2-8B 5 34.6 23.1 28.7 48.2 39.8 36.7 30.7 29.9 39.6
InternVL2-40B 3 36.0 34.9 26.9 46.5 31.8 42.1 32.2 34.0 39.6
@@
LLaVA-NeXT-Video-7B 4 35.6 48.5 14.0 47.8 24.2 43.5 42.4 34.0 30.6
LLaVA-NeXT-Video-72B 1 40.9 48.9 22.8 57.4 35.3 42.4 36.7 35.0 48.6
@@
Figure 7: Performance comparisons between Vision Enabled (w/ video), Vision Disabled (w/o video) and Chance Level (Freq.). Enabled-Disabled indicates the gap between Vision Enabled and Vision Disabled, and Disabled-Chance betokens the gap between Vision Disabled and Chance Level (Freq.). Tasks are sorted by Enable-Disable for better understanding.
@@ 4.1 Evaluation Setup
Benchmark Models. We comprehensively evaluate 15 video-supporting MLLMs across diverse model families, encompassing various parameter scales and training recipes. For proprietary models, we consider Gemini-1.5 [76] and GPT-4o [33]. For open-source models, we evaluate models from InternVL2 [14], ViLA [44], LongViLA [88], LongVA [98], LLaVA-OneVision [39], and LLaVA-NeXT-Video [99]. All evaluations are conducted under zero-shot settings and using each model’s default prompts. To ensure reproducibility, we use greedy decoding for all models.
Metric Design. Based on whether the ground-truth answer is verbal or numerical, our tasks are suited to either a Multiple-Choice Answer (MCA) or Numerical Answer (NA) format (see Fig.˜3). For MCA tasks, we follow standard practice [96, 24, 30] by using Accuracy ($\mathcal{ACC}$), based on exact matching (with possible fuzzy matching), as the primary metric. For NA tasks, where models predict continuous values, accuracy via exact matching fails to capture the degree of proximity between model predictions and ground-truth answers. Therefore, we introduce a new metric, Mean Relative Accuracy ($\mathcal{MRA}$), inspired by previous works [22, 45, 71]. Specifically, for a NA question, given a model’s prediction $\hat{y}$, ground truth $y$, and a confidence threshold $\theta$, relative accuracy is calculated by considering $\hat{y}$ correct if the relative error rate, defined as $|\hat{y}-y|/y$, is less than $1-\theta$. As single-confidence-threshold accuracy only considers relative error in a narrow scope, $\mathcal{MRA}$ averages the relative accuracy across a range of confidence thresholds $\mathcal{C}=\{0.5,0.55,\dots,0.95\}$:
$\vskip-5.69046pt\small\mathcal{MRA}=\frac{1}{10}\sum_{\theta\in\mathcal{C}}\mathbbm{1}\left(\frac{|\hat{y}-y|}{y}<1-\theta\right).$ (1)
@@
Human Level Performance. We randomly sample a subset of 400 questions (50 per task), which we will refer to as VSI-Bench (tiny). Human evaluators independently answer each question, and their performance is evaluated using the above-mentioned metrics. For comparison, we also report Gemini-1.5 Pro’s performance on VSI-Bench (tiny). See Appendix˜C for details on evaluation setups.
@@ 4.2 Main Results
Fig.˜6 shows overall model performance on VSI-Bench. Our key observations are as follows:
@@
Open-source MLLMs. Top-tier open-source models like LLaVA-NeXT-Video-72B and LLaVA-OneVision-72B demonstrate highly competitive performance to closed-source models, trailing the leading Gemini-1.5 Pro by only 4% to 5%. However, the majority of open-source models (7//12) perform below the chance level baseline, indicating significant limitations in their visual-spatial intelligence.
Blind Evaluation. We compare MLLMs’ performance against “Chance Level (frequency)” and “Vision Disabled” (blind) results, using averages across six of the strongest models (3 open-source and 3 closed-source). As shown in Fig.˜7, the consistent improvements in “Enabled-Disabled” and general degradation in “Disabled-Chance” demonstrate that video is essential and beneficial for our VSI-Bench, with blind models performing below chance level. Meanwhile, MLLMs struggle to improve beyond chance level in the absolute distance estimation, route plan, and relative direction tasks, whether vision is enabled or not, underscoring the difficulty of these tasks. Note that on object size, “Vision Disabled” models already significantly outperform chance level, likely due to common-sense knowledge learned during language model training.
@@ 5 How MLLMs Think in Space Linguistically
To better understand when and why models succeed or fail and to elucidate the facets of visual-spatial intelligence they possess, we examine how MLLMs think in space linguistically here and visually in Sec.˜6. We begin by prompting the best-performing MLLM in VSI-Bench, Gemini-1.5 Pro [76], to articulate its internal reasoning in language.
@@ 5.1 Probing via Self-Explanations
Case Studies. Fig.˜8 presents self-explanations in both a success and an error case. In both examples, when thinking in space, the MLLM exhibits advanced video understanding, demonstrated by the impressive accuracy of its timestamped descriptions. The model also forms correct step-by-step reasoning processes, outlining steps such as “orient yourself”, “locate the dishwasher” and “visualize the quadrants” for the relative direction task. Furthermore, the construction of a global coordinate system (Fig.˜8, left) suggests that MLLMs may possess or build an implicit world model. Rather than using isolated frames, short clips, or random guesses, the MLLM used global spatial context and reasoning to infer correctly.
In the incorrect example (Fig.˜8, right), we can identify faulty visual-spatial capabilities like egocentric-allocentric transformation and relational reasoning, as introduced in Fig.˜2. In the video, the camera pans right to shift the view from the edge of the bed to the wall and window. The model obeys this egocentric view, responding that “to face the wall where the window is located, you must turn right” instead of creating an allocentric view reflecting the reality that the route from door to bed means turning left.
@@
Error Analysis. To quantify and identify the main bottleneck for the best-performing MLLM on our benchmark, we analyze its errors on VSI-Bench (tiny), categorizing them into four distinct types which arose from both our outlined visual-spatial capabilities (Fig.˜2) and a clear four-way categorization of errors upon examination:
@@
As shown in Fig.˜9, around 71% of errors are attributed to spatial reasoning (as ontologically conceived in Fig.˜2), which suggests that:
@@
Further analysis and case studies are in Sec.˜E.2.
@@ 5.2 Limits of CoT Methods in Visuospatial Tasks
Prompting techniques improve the reasoning and problem-solving abilities for large models across diverse tasks [34, 82, 32, 73]. Their successes motivate us to investigate whether these linguistic prompting methods could also improve the visual-spatial capabilities of MLLMs in VSI-Bench. We investigate three prevailing prompting techniques (see Sec.˜B.3 for more details):
@@
As shown in Fig.˜10, surprisingly, all three linguistic reasoning techniques lead to performance degradation on VSI-Bench. Zero-Shot CoT and ToT reduce average performance by about 4%, and self-consistency, though slightly better, still falls 1.1% below the no-prompting baseline. The unilateral improvement in the appearance order and absolute distance estimation tasks is easily explained by their significant percentage of linguistic intelligence errors (see Fig.˜9). In contrast, the room size and object size tasks suffer a large 8% to 21% decrease, showing that encouraging a model to think more can be not just unreliable but downright harmful. Meanwhile, as shown in Tab.˜1, ZeroShot CoT achieves a 1.6% improvement on the general video understanding benchmark VideoMME [24]. Therefore, our results suggest that:
@@ 6.1 Probing via Cognitive Maps
We prompt MLLMs to express their internal representations of the spaces they see using cognitive maps, a well-established framework for remembering objects in a set environment [60, 78]. We prompt the best-performing MLLM, Gemini-1.5 Pro, to predict object center positions within a 10 $\times$ 10 grid based on video input (see Fig.˜12b for grid size ablation and Sec.˜B.4 for prompt). We present examples of the resulting cognitive maps in Fig.˜11.
To quantitatively assess these cognitive maps, we evaluate the Euclidean distance between all pairs of objects within each map. We consider the distance (on the grid) between two objects to be correct if it deviates by no more than one grid unit from the distance in the ground truth cognitive map. As shown in Fig.˜12, we divide the map-distances into eight distinct bins for analysis. Interestingly, we find that the MLLM achieves a remarkable 64% accuracy in positioning adjacent objects within its cognitive map, indicating robust local spatial awareness. However, this accuracy significantly deteriorates as the distance between two objects increases, which suggests that:
@@
Figure 11: Visualization of cognitive maps from MLLM and GT.
@@ 6.2 Better Distance Reasoning via Cognitive Maps
Given the local awareness of MLLMs in remembering spaces (see Fig.˜11 and Fig.˜12) and the importance of mental imagery to how humans think in space, we investigate whether generating and using cognitive maps can help MLLMs’ spatial reasoning in terms of VSI-Bench’s relative distance task. This tests if the local distance awareness emerged through cognitive maps transfers to improved distance recall and reasoning.
@@
We prompt Gemini-1.5 Pro to first generate a cognitive map based on the given video and question, and then to use the predicted map to answer the question. As shown in Tab.˜2a, we find that using mental imagery improves a MLLM’s relative distance accuracy by 10%. The 20% to 32% gain over baseline with the ground truth cognitive map underscores the importance of building accurate mental maps of a scene, which enforce a globally consistent topology, but indicates that such mental imagery is only one part of the puzzle, albeit a crucial one. These results point to building a mental spatial world model or cognitive map as a valuable pretext task or a promising solution for MLLMs to tackle visual-spatial reasoning.

### v2

@@ 1 Introduction
Visual-spatial intelligence entails perceiving and mentally manipulating spatial relationships [26]; it requires myriad capabilities, including relational reasoning and the ability to transform between egocentric and allocentric perspectives (Sec. 2). While Large Language Models (LLMs) [61, 88, 6, 105, 36, 67, 68, 9, 81, 82, 3, 77] have advanced linguistic intelligence, visual-spatial intelligence remains under-explored, despite its relevance to robotics [21, 8, 7, 64], autonomous driving [79], and AR/VR [12, 27, 55].
Multimodal Large Language Models (MLLMs) [34, 78, 49, 1, 42, 49, 4, 15], which integrate language and vision, exhibit strong capacities to think and reason in open-ended dialog and practical tasks like web agents [35, 28, 33, 21]. To advance this intelligence in the visual-spatial realm, we introduce VSI-Bench, a video-based benchmark featuring over 5,000 question-answer pairs across nearly 290 real indoor-scene videos (Sec. 3). Video data, by capturing continuous, temporal input, both parallels how we observe the world and enables richer spatial understanding and reasoning than static images. Evaluating open- and closed-source models on VSI-Bench reveals that even though a large performance gap exists between models and humans, MLLMs exhibit emerging visual-spatial intelligence despite the challenges of video understanding, textual understanding, and spatial reasoning (Sec. 4).
To analyze model behavior and inspired by dual-coding theory [18], which posits that linguistic and visual processing are distinct yet complementary, we prompt selected models for self-explanations (linguistic) and cognitive maps (visual). Analyzing the self-explanations reveals that spatial reasoning, as compared to visual perception, linguistic intelligence, or temporal processing, is the main factor behind weak performance on VSI-Bench (Sec. 5). “Cognitive maps”, which represent internal layouts of environments [80, 62], allow us to evaluate MLLMs’ implicit spatial world models and find that MLLMs build strong local models but weak global ones (Sec. 6). Furthermore, standard linguistic reasoning techniques fail to enhance performance on our benchmark. However, explicitly generating and using cognitive maps improves spatial distance question-answering.
@@ 2 Visual-Spatial Intelligence
Taxonomy. We provide a taxonomy of capabilities potentially required for visual-spatial intelligence (Fig. 2), based on cognitive psychology [26, 57, 62, 11] and human experience with our benchmark tasks in Sec. 3. Visual perception, linguistic intelligence, temporal processing, and spatial reasoning are the four areas needed in VSI-Bench. For example, [11] shows that visual object and spatial processing are neurally distinct, which motivates “visual perception” and “spatial reasoning” as separate areas. We break spatial reasoning into two broad capabilities: relational reasoning and egocentric-allocentric transformation.
@@ 3.1 Overview
VSI-Bench includes eight tasks of three types: configurational, measurement estimation, and spatiotemporal. The configurational tasks (object count, relative distance, relative direction, route plan) test a model’s understanding of the configuration of a space and are more intuitive for humans (see Sec. 4 for comparison between MLLM and human performance). Measurement estimation (of object size, room size, and absolute distance) is of value to any embodied agent. While predicting a measurement exactly is very difficult, for both humans and models, a better sense of distance and other measurements is intuitively correlated with better visual-spatial intelligence and underpins a wide range of tasks that require spatial awareness, like interaction with objects and navigation. Spatiotemporal tasks like appearance order test a model’s memory of a space as seen in video. See Fig. 3 for an overview of VSI-Bench tasks and Fig. 5 for dataset statistics.
@@ 3.2 Benchmark Construction
We develop a sophisticated benchmark construction pipeline to effectively generate high-quality question-answer (QA) pairs at scale, as shown in Fig. 4.
@@
Question-Answer Generation. QA pairs are primarily auto-annotated using the meta-information and question templates; the route plan task was human-annotated. We sophisticatedly design and refine the question template for each task and provide guidelines for human annotators. For more detailed design, see Sec. B.1.
@@
InternVL2-2B 11 26.5 25.7 24.0 20.0 29.2 32.1 44.1 30.4 6.3
InternVL2-8B 3 37.5 31.3 29.0 48.9 44.2 38.0 33.4 28.9 46.4
InternVL2-40B 4 37.0 41.3 26.2 48.2 27.5 47.6 32.7 27.8 44.7
@@
LLaVA-Video-7B 5 35.6 48.5 14.0 47.8 24.2 43.5 42.4 34.0 30.6
LLaVA-Video-72B 1 40.9 48.9 22.8 57.4 35.3 42.4 36.7 35.0 48.6
@@ 4.1 Evaluation Setup
Benchmark Models. We comprehensively evaluate 15 video-supporting MLLMs across diverse model families, encompassing various parameter scales and training recipes. For proprietary models, we consider Gemini-1.5 [78] and GPT-4o [34]. For open-source models, we evaluate models from InternVL2 [14], ViLA [45], LongViLA [91], LongVA [101], LLaVA-OneVision [40], and LLaVA-Video [104]. All evaluations are conducted under zero-shot settings and using each model’s default prompts. To ensure reproducibility, we use greedy decoding for all models.
Metric Design. Based on whether the ground-truth answer is verbal or numerical, our tasks are suited to either a Multiple-Choice Answer (MCA) or Numerical Answer (NA) format (see Fig. 3). For MCA tasks, we follow standard practice [99, 24, 31] by using Accuracy ($\mathcal{ACC}$), based on exact matching (with possible fuzzy matching), as the primary metric. For NA tasks, where models predict continuous values, accuracy via exact matching fails to capture the degree of proximity between model predictions and ground-truth answers. Therefore, we introduce a new metric, Mean Relative Accuracy ($\mathcal{MRA}$), inspired by previous works [22, 46, 73]. Specifically, for a NA question, given a model’s prediction $\hat{y}$, ground truth $y$, and a confidence threshold $\theta$, relative accuracy is calculated by considering $\hat{y}$ correct if the relative error rate, defined as $|\hat{y}-y|/y$, is less than $1-\theta$. As single-confidence-threshold accuracy only considers relative error in a narrow scope, $\mathcal{MRA}$ averages the relative accuracy across a range of confidence thresholds $\mathcal{C}=\{0.5,0.55,\dots,0.95\}$:
$\vspace{-0.2cm}\small\mathcal{MRA}=\frac{1}{10}\sum_{\theta\in\mathcal{C}}\mathbbm{1}\left(\frac{|\hat{y}-y|}{y}<1-\theta\right).$ (1)
@@
Human Level Performance. We sample a subset of 400 questions (50 per task), which we will refer to as VSI-Bench (tiny). Human evaluators independently answer each question, and their performance is evaluated using the above-mentioned metrics. For comparison, we also report Gemini-1.5 Pro’s performance on VSI-Bench (tiny). See Appendix C for details on evaluation setups.
@@ 4.2 Main Results
Fig. 6 shows overall model performance on VSI-Bench. Our key observations are as follows:
@@
Open-source MLLMs. Top-tier open-source models like LLaVA-Video-72B and LLaVA-OneVision-72B demonstrate highly competitive performance to closed-source models, trailing the leading Gemini-1.5 Pro by only 4% to 5%. However, the majority of open-source models (7///12) perform below the chance level baseline, indicating significant limitations in their visual-spatial intelligence.
@@ 5 How MLLMs Think in Space Linguistically
To better understand when and why models succeed or fail and to elucidate the facets of visual-spatial intelligence they possess, we examine how MLLMs think in space linguistically here and visually in Sec. 6. We begin by prompting the best-performing MLLM in VSI-Bench, Gemini-1.5 Pro [78], to articulate its internal reasoning in language.
@@ 5.1 Probing via Self-Explanations
Case Studies. Fig. 7 presents self-explanations in both a success and an error case. In both examples, when thinking in space, the MLLM exhibits advanced video understanding, demonstrated by the impressive accuracy of its timestamped descriptions. The model also forms correct step-by-step reasoning processes, outlining steps such as “orient yourself”, “locate the dishwasher” and “visualize the quadrants” for the relative direction task. Furthermore, the construction of a global coordinate system (Fig. 7, left) suggests that MLLMs may possess or build an implicit world model. Rather than using isolated frames, short clips, or random guesses, the MLLM used global spatial context and reasoning to infer correctly.
@@
In the incorrect example (Fig. 7, right), we can identify faulty visual-spatial capabilities like egocentric-allocentric transformation and relational reasoning, as introduced in Fig. 2. In the video, the camera pans right to shift the view from the edge of the bed to the wall and window. The model obeys this egocentric view, responding that “to face the wall where the window is located, you must turn right” instead of creating an allocentric view reflecting the reality that the route from door to bed means turning left.
Error Analysis. To quantify and identify the main bottleneck for the best-performing MLLM on our benchmark, we analyze its errors on VSI-Bench (tiny), categorizing them into four distinct types which arose from both our outlined visual-spatial capabilities (Fig. 2) and a clear four-way categorization of errors upon examination:
@@
As shown in Fig. 8, around 71% of errors are attributed to spatial reasoning (as ontologically conceived in Fig. 2), which suggests that:
@@
Further analysis and case studies are in Sec. E.2.
@@ 5.2 Limits of CoT Methods in Visuospatial Tasks
Prompting techniques improve the reasoning and problem-solving abilities for large models across diverse tasks [35, 84, 33, 75]. Their successes motivate us to investigate whether these linguistic prompting methods could also improve the visual-spatial capabilities of MLLMs in VSI-Bench. We investigate three prevailing prompting techniques (see Sec. B.3 for more details):
@@
As shown in Fig. 9, surprisingly, all three linguistic reasoning techniques lead to performance degradation on VSI-Bench. Zero-Shot CoT and ToT reduce average performance by about 4%, and self-consistency, though slightly better, still falls 1.1% below the no-prompting baseline. The unilateral improvement in the appearance order and absolute distance estimation tasks is easily explained by their significant percentage of linguistic intelligence errors (see Fig. 8). In contrast, the room size and object size tasks suffer a large 8% to 21% decrease, showing that encouraging a model to think more can be not just unreliable but downright harmful. Meanwhile, as shown in Tab. 1, ZeroShot CoT achieves a 1.6% improvement on the general video understanding benchmark VideoMME [24]. Therefore, our results suggest that:
@@ 6.1 Probing via Cognitive Maps
We prompt MLLMs to express their internal representations of the spaces they see using cognitive maps, a well-established framework for remembering objects in a set environment [62, 80]. We prompt the best-performing MLLM, Gemini-1.5 Pro, to predict object center positions within a 10 $\times$ 10 grid based on video input (see Fig. 11b for grid size ablation and Sec. B.4 for prompt). We present examples of the resulting cognitive maps in Fig. 10.
Figure 10: Visualization of cognitive maps from MLLM and GT.
To quantitatively assess these cognitive maps, we evaluate the Euclidean distance between all pairs of objects within each map. We consider the distance (on the grid) between two objects to be correct if it deviates by no more than one grid unit from the distance in the ground truth cognitive map. As shown in Fig. 11, we divide the map-distances into eight distinct bins for analysis. Interestingly, we find that the MLLM achieves a remarkable 64% accuracy in positioning adjacent objects within its cognitive map, indicating robust local spatial awareness. However, this accuracy significantly deteriorates as the distance between two objects increases, which suggests that:
@@ 6.2 Better Distance Reasoning via Cognitive Maps
Given the local awareness of MLLMs in remembering spaces (see Fig. 10 and Fig. 11) and the importance of mental imagery to how humans think in space, we investigate whether generating and using cognitive maps can help MLLMs’ spatial reasoning in terms of VSI-Bench’s relative distance task. This tests if the local distance awareness emerged through cognitive maps transfers to improved distance recall and reasoning.
@@
We prompt Gemini-1.5 Pro to first generate a cognitive map based on the given video and question, and then to use the predicted map to answer the question. As shown in Tab. 2a, we find that using mental imagery improves a MLLM’s relative distance accuracy by 10%. The 20% to 32% gain over baseline with the ground truth cognitive map underscores the importance of building accurate mental maps of a scene, which enforce a globally consistent topology, but indicates that such mental imagery is only one part of the puzzle, albeit a crucial one. These results point to building a mental spatial world model or cognitive map as a valuable pretext task or a promising solution for MLLMs to tackle visual-spatial reasoning.